### v1

@@ 5.1.1 Stealth–Success Tradeoff Visualization
Despite similar degradation ($\sim 25\text{--}26$ dB), we see that Cheng2020-Attention is most robust (stealth PSNR $\sim 49$ dB); Cheng2020-Anchor achieves the same degradation with smaller visible perturbation (stealth PSNR $\sim 56$ dB); TCM is comparatively vulnerable (stealth PSNR $\sim 54$ dB), indicating susceptibility of its hybrid architecture.
@@ 5.2 Qualitative Analysis
The additive attack on Figure 5(f) introduces widespread high-frequency artifacts across all wavelet subbands, while Pixel-LogExp on Figure 5(g) shows reduced but still substantial spectral noise. In contrast, our T-MLA method in Figure 5(h) produces minimal frequency-domain disturbance, a critical advantage for evading spectral anomaly detection systems. This difference is particularly evident in the magnified frequency insets: while baseline methods show obvious perturbation patterns even to the naked eye, T-MLA perturbations are only discernible in the red inset, with the blue inset region appearing nearly identical to the original.
@@ 6 Conclusion
We introduced T-MLA, a targeted multiscale wavelet-domain attack that exploits vulnerabilities in neural image compression (NIC) models. By injecting frequency-specific perturbations, T-MLA achieves substantial degradation in reconstruction quality (25-30 dB PSNR drop) while preserving high imperceptibility (50-55 dB PSNR). Unlike pixel-space attacks, our method minimizes spectral artifacts, making detection more challenging. Across Kodak, CLIC, and DIV2K and three NIC models, attack impact is inversely related to image entropy: smoother images are more vulnerable. Weaker correlations on more diverse data (e.g., DIV2K) indicate additional complexity factors beyond texture. These results point to a broader vulnerability in VAE-based compression systems, which rely on entropy models that can be exploited through latent space perturbations. Future work includes improving model robustness, studying black-box and universal attacks, and extending our approach to other VAE-based systems such as video codecs, NeRFs, and diffusion models. Our findings raise concerns about the security of generative compression pipelines and their broader deployment in modern AI systems.
@@
All datasets used in this study are publicly available (Kodak, CLIC-pro, DIV2K). Source code and experiment scripts to reproduce the results are available at https://github.com/nkalmykovsk/tmla (commit hash and instructions are provided in the repository). No proprietary data were used.
@@
We evaluate T-MLA in the white-box setting on three NICs; black-box/transfer variants and broader codec coverage (e.g., mbt2018) are deferred to future work. Appendix F outlines an initial wavelet-aware defense; its effectiveness (including against PGD/LogExp) will be evaluated and quantified in future work. Our goal is to provide a principled multiscale wavelet-domain framework that couples stealth and degradation, enabling reproducible robustness studies.
@@ B.1 Entropy Calculation
We quantify image complexity using normalized local entropy. Given an RGB image $\mathbf{x}\in\mathbb{R}^{M\times N}$ with $I\in[0,1]^{I\times J\times 3}$, we convert it to grayscale $Y\in\{0,\ldots,255\}^{I\times J}$ and compute local Shannon entropy in a circular neighborhood $\mathcal{N}_{r}(x,y)$ of radius $r=10$ (empirically balances locality and statistical stability):
@@
$\mu_{E}=\frac{1}{I\cdot J}\sum_{x=1}^{I}\sum_{y=1}^{J}\tilde{E}(x,y).$
@@ B.2 Full Correlation Analysis
Figure B.3 shows correlation analysis between mean entropy and relative VIF drop across all datasets and compression models. For DIV2K and CLIC datasets, we excluded $10\%$ of samples with mean entropy <0.6<0.6 and relative VIF drop <0.1<0.1 due to significant variations in depth maps and texture distributions across these images (see in Figure G.7, Figure G.8, and Figure G.9).
@@
Dataset variations reveal factors beyond entropy affecting model vulnerability: Kodak shows the strongest correlations (r=0.579r=-0.579), while DIV2K and CLIC show weaker correlations (r=0.241r=-0.241). These differences stem from dataset characteristics, Kodak contains professionally captured photographs with controlled lighting and predictable texture distributions, while DIV2K and CLIC feature more diverse natural images and compression-optimized content, respectively.
@@ C.5 Relation to our attack
Natural images exhibit rapidly decaying energy in the detail bands, most coefficients are near zero. Our log–exponential perturbation strategy injects noise into ten subbands simultaneously (all detail bands and only the final approximation band $LL^{(3)}$ due to the pyramidal dependency), deliberately deviating from this sparsity pattern. This targeted high-frequency perturbation approach makes the noise particularly difficult for the human eyes to detect, while maximizing downstream distortion.
@@ Appendix F Defense Studies
Figure F.6: Defended image for DIV525 dataset.
@@ F.1 Algorithm Overview
Appendix G Additional Results
In this section, we provide detailed quantitative results in Table G.4, Table G.5, Table G.7, Table G.8, Table G.9, Table G.11, Table G.12, Table G.13, Table G.15, Table G.17, Table G.19. These tables present comprehensive per-image attack performance metrics for three models (Cheng2020-Anchor, Cheng2020-Attention, and TCM) evaluated on Kodak, CLIC, and DIV2K datasets. For each image, we report: stealth metrics ($\uparrow$): PSNR and VIF scores measuring imperceptibility; attack success metrics ($\downarrow$): PSNR, SSIM, and VIF measuring distortion; compression rate in bits-per-pixel (BPP).
Figure G.7: Example of DIV657 dataset.
Figure G.8: Example of DIV155 dataset.
Figure G.9: Example of DIV049 dataset.
Table G.4: Detailed evaluation results for Cheng2020-Anchor model on Kodak dataset
@@ Appendix G Additional Results
Table G.5: Detailed evaluation results for Cheng2020-Anchor model on CLIC subset
@@
Table G.6: Detailed evaluation results for Cheng2020-Anchor model on CLIC subset (continued)
@@
Table G.7: Detailed evaluation results for Cheng2020-Anchor model on DIV2K subset
@@
Table G.8: Detailed evaluation results for Cheng2020-Anchor model on DIV2K subset (continued)
@@
Table G.9: Detailed evaluation results for Cheng2020-Attention model on Kodak dataset
@@
Table G.10: Detailed evaluation results for Cheng2020-Attention model on CLIC subset
@@
Table G.11: Detailed evaluation results for Cheng2020-Attention model on CLIC subset (continued)
@@
Table G.12: Detailed evaluation results for Cheng2020-Attention model on DIV2K subset
@@
Table G.13: Detailed evaluation results for Cheng2020-Attention model on DIV2K subset (continued)
@@
Table G.14: Detailed evaluation results for LIC-TCM model on Kodak dataset
@@
Table G.15: Detailed evaluation results for LIC-TCM model on Kodak dataset (continued)
@@
Table G.16: Detailed evaluation results for LIC-TCM model on CLIC subset
@@
Table G.17: Detailed evaluation results for LIC-TCM model on CLIC subset (continued)
@@
Table G.18: Detailed evaluation results for LIC-TCM model on DIV2K subset
@@
Table G.19: Detailed evaluation results for LIC-TCM model on DIV2K subset (continued)

### v2

@@ 5.1.1 Stealth–Success Tradeoff Visualization
Despite similar degradation ($\sim 25\text{--}26$ dB), we see that Cheng2020-Attention is most robust (stealth PSNR $\sim 49$ dB); Cheng2020-Anchor achieves the same degradation with smaller visible perturbation (stealth PSNR $\sim 56$ dB); LIC-TCM is comparatively vulnerable (stealth PSNR $\sim 54$ dB), indicating susceptibility of its hybrid architecture.
@@
5.1.3 Optimization Convergence and Efficiency
We next analyze the convergence behavior and computational cost of T-MLA. A detailed convergence analysis, including $Q_{\text{in}}$ and $Q_{\text{out}}$ versus iterations for a representative case (kodak23, Cheng2020-Anchor), is provided in Appendix G. The plots show that the input PSNR is increased in stages towards the 55 dB target while the output PSNR remains concentrated around 25 dB, indicating stable optimization under the joint stealth–degradation objective. Precise runtimes and GPU memory usage for all three NIC models are summarized in section 4.
5.1.4 Perceptual Metrics Analysis
[…]
@@ 5.2 Qualitative Analysis
The additive attack in Figure 5(f) introduces widespread high-frequency artifacts across all wavelet subbands, while Pixel-LogExp in Figure 5(g) shows reduced but still substantial spectral noise. In contrast, our T-MLA method in Figure 5(h) produces minimal frequency-domain disturbance, a critical advantage for evading spectral anomaly detection systems. This difference is particularly evident in the magnified frequency insets: while baseline methods show obvious perturbation patterns even to the naked eye, T-MLA perturbations are only discernible in the red inset, with the blue inset region appearing nearly identical to the original.
@@
5.3 Defense Evaluation
We explored a wavelet-aware defense mechanism to counter T-MLA, conceptually mirroring the attack by minimizing distortion via gradient descent in the wavelet domain (detailed in Appendix F). While a full benchmark of defense strategies (e.g., adversarial training, frequency anomaly detection) is beyond our scope, preliminary results (see Figure F.6) show that this adaptive defense can effectively restore visual quality (PSNR/SSIM) by neutralizing adversarial perturbations in the frequency bands they target. This confirms that understanding wavelet-domain vulnerabilities is key to designing robust NIC systems.
@@ 6 Conclusion
We introduced T-MLA, a targeted multiscale wavelet-domain attack that exploits vulnerabilities in neural image compression (NIC) models. By injecting frequency-specific perturbations, T-MLA achieves substantial degradation in reconstruction quality (25-30 dB PSNR drop) while preserving high imperceptibility (50-55 dB PSNR). Unlike pixel-space attacks, our method minimizes spectral artifacts, making detection more challenging. Across the Kodak, CLIC, and DIV2K datasets and three NIC models, attack impact is inversely related to image entropy: smoother images are more vulnerable. Weaker correlations on more diverse data (e.g., DIV2K) indicate additional complexity factors beyond texture. These results point to a broader vulnerability in VAE-based compression systems, which rely on entropy models that can be exploited through latent space perturbations.
From a practical standpoint, our findings suggest that NIC deployment in safety-critical or high-value pipelines (e.g., medical imaging, autonomous driving, commercial CDNs) requires new defense layers. Specifically, we recommend integrating wavelet-domain anomaly detection to flag high-frequency spectral irregularities and adopting entropy-adaptive coding schemes that allocate redundant bits to low-texture regions, thereby hardening the entropy bottleneck against targeted perturbations.
Future work includes improving model robustness, studying transfer and query-based black-box variants of T-MLA across heterogeneous NIC architectures, and extending our approach to other Variational Autoencoder (VAE)-based systems such as video codecs, Neural Radiance Fields (NeRFs), and diffusion models. Our findings raise concerns about the security of generative compression pipelines and their broader deployment in modern AI systems.
@@
All datasets used in this study are publicly available (Kodak, CLIC-pro, DIV2K). Source code and experiment scripts to reproduce the results are available at https://github.com/nkalmykovsk/tmla (instructions are provided in the repository). No proprietary data were used.
@@
We evaluate T-MLA in the white-box setting on three NICs; black-box/transfer variants and broader codec coverage (e.g., mbt2018) are deferred to future work. Appendix F outlines an initial wavelet-aware defense. Our goal is to provide a principled multiscale wavelet-domain framework that couples stealth and degradation, enabling reproducible robustness studies.
@@ B.1 Entropy Calculation
We quantify image complexity using normalized local entropy. Given an RGB image $\mathbf{x}\in[0,1]^{H\times W\times 3}$, we convert it to grayscale $Y\in\{0,\ldots,255\}^{H\times W}$ and compute local Shannon entropy in a circular neighborhood $\mathcal{N}_{r}(x,y)$ of radius $r=10$ (empirically balances locality and statistical stability):
@@
$\mu_{E}=\frac{1}{H\cdot W}\sum_{x=1}^{H}\sum_{y=1}^{W}\tilde{E}(x,y).$
@@ B.2 Full Correlation Analysis
Figure B.3 shows correlation analysis between mean entropy and relative VIF drop across all datasets and compression models. For DIV2K and CLIC datasets, we excluded $10\%$ of samples with mean entropy <0.6<0.6 and relative VIF drop <0.1<0.1. This filtering removes outliers with significant variations in depth maps and non-standard texture distributions (see examples in Figure I.10, Figure I.11, and Figure I.12). Crucially, this selection ensures that the complexity distribution of the larger datasets aligns with the standard Kodak benchmark, preventing results from being skewed by trivial, low-information samples and ensuring generalization to realistic, detail-rich content.
@@
Dataset variations reveal factors beyond entropy affecting model vulnerability: Kodak shows the strongest correlations (r=0.579r=-0.579), while DIV2K and CLIC show weaker correlations (r=0.241r=-0.241). These differences stem from dataset characteristics: Kodak contains professionally captured photographs with controlled lighting and predictable texture distributions, while DIV2K and CLIC feature more diverse natural images and compression-optimized content, respectively.
@@ C.5 Relation to our attack
Natural images exhibit rapidly decaying energy in the detail bands; most coefficients are near zero. Our log–exponential perturbation strategy injects noise into ten subbands simultaneously (all detail bands and only the final approximation band $LL^{(3)}$ due to the pyramidal dependency), deliberately deviating from this sparsity pattern. This targeted high-frequency perturbation approach makes the noise particularly difficult for the human eyes to detect, while maximizing downstream distortion.
@@ Appendix F Defense Studies
Figure F.6: Defended image for DIV090 image. (a) The adversarial perturbed image ($\mathrm{PSNR}=55.0$ dB, $\mathrm{VIF}=0.999$) looks clean to the human eye. (b) The attacked reconstruction ($\mathrm{PSNR}=24.2$ dB, $\mathrm{VIF}=0.828$) exhibits visible artifacts, particularly on the player’s arm. (c) The defended image ($\mathrm{PSNR}=39.2$ dB, $\mathrm{VIF}=0.999$) obtained via our iterative defense (Algorithm 2) restores quality: while PSNR is lower than the stealthy input, it is higher than the attacked output, and artifacts are visually removed.
@@ F.1 Algorithm Overview
Appendix G Convergence and Sensitivity Analysis
We analyze the convergence behavior of T-MLA on a representative case (kodak23, Cheng2020-Anchor, S=3S{=}3, $Q_{\text{in}}{=}55$ dB, $Q_{\text{out}}{=}25$ dB). As shown in Figure G.7, the input PSNR $Q_{\text{in}}$ increases in stages: the optimizer reaches an intermediate PSNR level, briefly stabilizes, and then moves to the next, tighter constraint until converging near 55 dB. Throughout optimization, the output PSNR $Q_{\text{out}}$ remains concentrated around the 25 dB target, indicating that the joint stealth–degradation objective is enforced reliably.
Regarding hyperparameters, we empirically select the largest feasible $Q_{\text{in}}$ and the smallest feasible $Q_{\text{out}}$ for which optimization converges. Very low $Q_{\text{out}}$ targets (extreme degradation) or excessively large scale factors $\alpha$ shrink the feasible perturbation region and can destabilize convergence; in practice, small changes of $\alpha$ (e.g., $\pm 0.1$ around $1.8$) do not noticeably affect convergence. In the main experiments we fix $(Q_{\text{in}},Q_{\text{out}})=(50\text{--}55,25)$ dB, $\alpha=1.8$, and a moderate sparsity weight $\lambda$; in contrast to $\alpha$, the method is sensitive to $\lambda$, and even changes at the third decimal place can alter convergence behavior, with too large $\lambda$ enforcing overly sparse perturbations and degrading convergence.
(a) Input PSNR $Q_{\text{in}}$.
(b) Output PSNR $Q_{\text{out}}$.
Figure G.7: Convergence of T-MLA on Kodak23 with Cheng2020-Anchor (S=3S{=}3 scales, $Q_{\text{in}}{=}55$ dB, $Q_{\text{out}}{=}25$ dB). The input PSNR (left) progressively increases towards the target while the output PSNR (right) remains constrained near 25 dB, illustrating stable optimization under the joint stealth–degradation objective.
Appendix H Perceptual Robustness Evaluation
Figure H.8: Perceptual robustness evaluation on Cheng2020-Anchor across Kodak dataset. (a) Trade-off between LPIPS (log scale) and PSNR: T-MLA (blue crosses) stays well below the visibility threshold ($0.01$) with high PSNR, unlike PGD (red circles). (b) VIF score distribution: T-MLA maintains values near 1.0 (best). (c) DISTS score distribution: T-MLA yields scores near 0 (best), indicating minimal texture distortion compared to PGD.
To validate the perceptual invisibility of our adversarial perturbations, we extended our evaluation beyond standard MSE-based metrics (PSNR, SSIM) to include deep learning-based perceptual metrics: LPIPS (Learned Perceptual Image Patch Similarity) and DISTS (Deep Image Structure and Texture Similarity).
[…]
While we observe that PSNR and LPIPS do not perfectly correlate in the high-quality regime ($PSNR>50$ dB) due to the sensitivity of deep metrics to microscopic noise, the global trend is consistent: the massive PSNR gain (+33.3 dB over baseline) translates directly into imperceptible LPIPS/DISTS scores, confirming that our perturbations are perceptually invisible.
Appendix I Additional Results
In this section, we provide detailed quantitative results in Table I.5, Table I.6, Table I.8, Table I.9, Table I.10, Table I.12, Table I.13, Table I.14, Table I.16, Table I.18, Table I.20. These tables present comprehensive per-image attack performance metrics for three models (Cheng2020-Anchor, Cheng2020-Attention, and LIC-TCM) evaluated on Kodak, CLIC, and DIV2K datasets. For each image, we report: stealth metrics ($\uparrow$): PSNR and VIF scores measuring imperceptibility; attack success metrics ($\downarrow$): PSNR, SSIM, and VIF measuring distortion; compression rate in bits-per-pixel (BPP).
Figure I.9: Statistical comparison between the baseline PGD attack and our T-MLA on the Kodak dataset (24 images). Bars show mean $\pm$ standard deviation of Stealth PSNR (left, higher is better) and Attack VIF (right, lower is better) for Cheng2020-Anchor, Cheng2020-Attention, and LIC-TCM. All improvements of T-MLA over PGD are statistically significant according to a two-sided paired t-test on per-image metrics ($p<10^{-8}$ for all cases).
Table I.4: Summary statistics of different attack methods for Cheng2020-Anchor on the Kodak dataset (24 images).
Method Stealth $\uparrow$ Attack Success $\downarrow$ PSNR VIF PSNR SSIM VIF BPP PGD $40.15\pm 0.02$ $0.973\pm 0.018$ $8.89\pm 2.73$ $691\pm 0.103 0.0$$081\pm 0.137$ $0.93\pm 0.37$ APGD [croce2020apgd] $42.10\pm 0.30$ $0.980\pm 0.017$ $25.18\pm 1.23$ $0.953\pm 0.022$ $0.618\pm 0.308$ $0.95\pm 0.36$ LogExp $43.42\pm 0.33$ $0.984\pm 0.004$ $21.30\pm 1.52$ $0.910\pm 0.090$ $0.488\pm 0.117$ $0.87\pm 0.35$ T-MLA $\mathbf{55.82\pm 0.87}$ $\mathbf{0.999\pm 0.001}$ $25.03\pm 0.33$ $0.969\pm 0.010$ $0.728\pm 0.140$ $0.85\pm 0.37$
Figure I.10: Example of DIV657 dataset.
Figure I.11: Example of DIV155 dataset.
Figure I.12: Example of DIV049 dataset.
Table I.5: Detailed evaluation results for Cheng2020-Anchor model on Kodak dataset
@@ Appendix I Additional Results
Table I.6: Detailed evaluation results for Cheng2020-Anchor model on CLIC subset
@@
Table I.7: Detailed evaluation results for Cheng2020-Anchor model on CLIC subset (continued)
@@
Table I.8: Detailed evaluation results for Cheng2020-Anchor model on DIV2K subset
@@
Table I.9: Detailed evaluation results for Cheng2020-Anchor model on DIV2K subset (continued)
@@
Table I.10: Detailed evaluation results for Cheng2020-Attention model on Kodak dataset
@@
Table I.11: Detailed evaluation results for Cheng2020-Attention model on CLIC subset
@@
Table I.12: Detailed evaluation results for Cheng2020-Attention model on CLIC subset (continued)
@@
Table I.13: Detailed evaluation results for Cheng2020-Attention model on DIV2K subset
@@
Table I.14: Detailed evaluation results for Cheng2020-Attention model on DIV2K subset (continued)
@@
Table I.15: Detailed evaluation results for LIC-TCM model on Kodak dataset
@@
Table I.16: Detailed evaluation results for LIC-TCM model on Kodak dataset (continued)
@@
Table I.17: Detailed evaluation results for LIC-TCM model on CLIC subset
@@
Table I.18: Detailed evaluation results for LIC-TCM model on CLIC subset (continued)
@@
Table I.19: Detailed evaluation results for LIC-TCM model on DIV2K subset
@@
Table I.20: Detailed evaluation results for LIC-TCM model on DIV2K subset (continued)